# Concurrent-Allocation Task Execution for Multi-Robot Path-Crossing-Minimal Navigation in Obstacle Environments

### v1

@@ VI-A Accommodating the CATE to Unicycle Dynamics
TABLE II: Second simulation: Quantitative evaluation metrics of the multi-robot navigation efficiency with 16 different robot-obstacle group settings.
@@ VI-B3 CAPT[15] + ORCA [57]
Figure 11: Nine illustrative examples of the “Arrow” formation using the proposed CATE algorithm (19) in the second simulation. In the detailed nine trials, the numbers of robots and obstacles are set to be the same, i.e., $N=11$ and $M=7$, and the size of the obstacle circle is randomly selected in the region of $r\in[1.7,4]$. The positions of the robots and obstacles are randomly selected within the space of $\{x\in[-5,~{}35],y\in[0,~{}25]\}$, but do not overlap each other or in the space of $\{x\in[10,~{}25],y\in[7,~{}18]\}$. The abbreviation, blue and yellow triangles, red circles, and dashed colored lines have the same meanings as those in Fig. 6.
Figure 12: Three failure examples using the previous three different baseline methods in the second simulation. (a) The deadlock failure in the $7$-robot-$7$-obstacle scenario when using the FOTE method. (b) The infeasible failure in the $9$-robot-$6$-obstacle scenario when using the DAPT+ORCA method. (c) The infeasible failure in the $11$-robot-$7$-obstacle scenario when using the CAPT+ORCA method.
@@ VI-C Simulations
Figure 13: Third simulation: Effectiveness of the CATE algorithm (19) in environments with moving obstacles. (a) T=$1.2$s. (b) T=$3.5$s. (c) T=$20$s. The light blue circles represent the initial positions of the obstacles. The abbreviation, blue and yellow triangles, red circles, and dashed colored lines have the same meanings as those in Fig. 6.
@@
In the second simulations (see Tables I-II and Figs. 7-12), to better compare the performance with alternative approaches in Section VI-B, we conduct extensive experiments of different robot-obstacle combinations and analyze the aggregated results. Precisely, as shown in Table I, we select 16 groups of different numbers of robots and obstacles (i.e., robots: $N=5,7,9,11$, obstacles $M=4,5,6,7$), and then conduct 10 trials for each group. In each trial, the size of the obstacle circle is randomly selected in the range of $r\in[1.7,4]$. The positions of the robots and obstacles are randomly chosen within the area of $\{x\in[-5,~{}35],y\in[0,~{}25]\}$, but do not overlap other robots, obstacles, or in the center space of $\{x\in[10,~{}25],y\in[7,~{}18]\}$.
Figure 16: (a) Multi-AMR working space consisting of three different obstacles in Robotic Research Center (RRC). (b) Two AMRs SCOUTMINI employed in the experiments and components.
Figure 17: (a) 3D point map of the working space in Fig. 16 (a) built by a popular LiDAR SLAM framework: Fast-LIO. (b) 2D octomap from the x-y plane projection of the original 3D octomap [58]. (c) Structure of the multi-AMR onboard navigation system, which consists of onboard SLAM, onboard navigation, and onboard execution, respectively.
@@
As shown on the left side of Table I and Fig. 7, the success rates of the proposed CATE algorithm (19) are $100\%$ all along, which verifies the feasibility advantages of the algorithm in the obstacle environments. The right side of Table I and Fig. 8 illustrates the mean (SD) and boxplots of the convergence time of the four approaches, where the convergence time of the proposed CATE algorithm (19) is the smallest among the ones in three baseline benchmarks. To quantitatively evaluate the multi-robot navigation efficiency, we further employ two metrics for path crossings and trajectory lengths in Table II. It is observed on the left side of Table II that if the number of obstacles and robots increases, the proposed CATE algorithm (19) not only guarantees fewer path crossings than three baseline methods but also maintains $100\%$ success rates all along. As shown on the right side of Table II, the proposed CATE algorithm (19) only requires the smallest trajectory length than the ones in three baseline methods, which thus indicates the highest efficiency of the CATE algorithm as well. Moreover, Figs. 9 and 10 present the detailed boxplots of the path-crossing and trajectory-length metrics across four approaches, demonstrating that the CATE algorithm (19) achieves the highest performance.
@@
Since the initial positions of robots and obstacles are randomly selected in each trial for each algorithm, one has that these positions are different at each trial. Meanwhile, as shown in Table I, there are considerable differences in the success rates between the CAPT+ORCA and the CATE algorithms (19), whereas we only count the path crossings in Table II when the MPCM navigation is successfully achieved. Therefore, although the CAPT+ORCA algorithm achieves fewer path crossings than the proposed CATE algorithm (19) in Table II, it is not suitable to compare such path crossings directly. Here, we provide a detailed explanation below.
@@
(i) When the number of robots and obstacles is relatively small (i.e., $N=5,7$ and $M=4,5$), the random selection of their positions results in a high probability that obstacles will not be positioned between the robots and the desired points. Then, such random placement causes the environment to degenerate into an almost obstacle-free one, which is more suitable for the CAPT+ORCA algorithm (this also explains why the success rate of the CAPT+ORCA algorithms is low since it is effective in simplistic environments). (ii) When the numbers of robots and obstacles become relatively large (i.e., $N=9,11$ and $M=6,7$), although the CAPT+ORCA algorithm achieves relatively few path crossings, its success rate is quite low (only 40%-70%, as indicated in Table I). In contrast, the proposed CATE algorithm (19) not only guarantees a 100% success rate (see Table I) but also maintains relatively low path crossings. However, there is a tradeoff between the success rate and the number of path crossings. To achieve the 100% success rate, the CATE algorithm 19 needs to successfully deal with all complicated environments where obstacles effectively obstruct the motion from the initial positions to the desired positions of robots. This results in slightly more path crossings than the CAPT+ORCA algorithm. It’s worth mentioning that the CATE algorithm (19) significantly outperforms the other three algorithms in terms of convergence time and trajectory length in Tables I and II. Therefore, the superiority of the CATE algorithm (19) is verified as well.
@@
Figure 18: First experiment: Comparing platoon formations forming and snapshots using different methods in Robotic Research Center, NTU. (a) FOTE. (b) DAPT + ORCA. (c) CAPT + ORCA. (d) Proposed CATE. The abbreviation, blue, magenta and yellow triangles, red circles, dashed colored lines have the same meanings as those in Figs. 6 and 15.
@@
In the fifth simulation (see Fig. 15), to show the feasibility of the CATE algorithm (19) for possible deadlocks, we compare our algorithm with the fixed-ordering strategy in a special case of eight robots passing through a narrow gap. In particular, we use two big disk obstacles centered at $\mathbb{x}_{1}^{o}=[6,2]^{\mbox{\tiny\sf T}},\mathbb{x}_{2}^{o}=[6,14]^{\mbox{\tiny\sf T}}$ respectively with the same radius $r_{1}^{o}=r_{2}^{o}=5.5$ to occupy most of the passing space and leave a narrow gap deliberately. The initial positions of robots and desired points are set satisfying A2, A3, A4. As shown in Fig. 15 (a), even though eight robots governed by the CATE algorithm (19) are congestive at the entrance of the narrow gap (magenta vehicles), they can get rid of the congestion and finally, form a platoon (yellow vehicles) by flexibly reallocating their desired points with distinct ordering sequences, which verifies the strong feasibility of the CATE algorithm (19). However, one observes in Fig. 15 (b) that eight robots governed by the fixed-ordering methods fail to pass the gap and are trapped in deadlocks (magenta vehicles). Moreover, Fig. 15 (c) illustrates the evolution of the reallocated ordering sequences $s[i],i\in\mathbb{Z}_{1}^{8},$ for robot $i,i\in\mathcal{V},$ satisfying $s[i]=k,\alpha_{i,k}=1$, which converges to $s[1]=4,s[2]=5,s[3]=6,s[4]=2,s[5]=1,s[6]=3,s[7]=8,s[8]=7$ such that $\lim_{t\rightarrow\infty}\mathbb{x}_{i}(t)-\mathbb{x}_{s[i]}^{d}(t)=0$. Note that the final subscripts $s[i],i\in\mathbb{Z}_{1}^{8}$ are consistent with ones in Fig. 15 (a).
@@ VI-D Experiments with Two AMRs
In this subsection, we employ two AMRs (SCOUTMINI [^1] SCOUTMINI: https://global.agilex.ai/products/scout-mini) to validate Theorem 1. As shown in Fig. 16 (a), we establish a multi-AMR workspace, which consists of three different obstacles of a static vehicle, a workbench, and six cones occupying the way. As shown in Fig. 16 (b), each AMR is $0.615$m in length and $0.58$m in width and equipped with an onboard computer: NVIDIA Jetson Xavier NX3222NVIDIA Jetson Xavier NX3: https://www.nvidia.com/en-us/autonomous-machines/embedded-systems/jetson-xavier-nx/ and a Livox Lidar: MID-360333Livox Lidar MID-360: https://www.livoxtech.com/mid-360 or AVIA [^4] Livox Lidar AVIA: https://www.livoxtech.com/avia. However, since the AMR is not equipped with any localization equipment (such as an on-board GPS or off-board motion capture cameras), we utilize the Lidar with the popular FAST-LIO SLAM algorithm [59] to build a 3D point map (see, Fig. 17 (a)) and localize itself. Then, we get a 2D map from the projection of octomap using octotree [58] (i.e., red circles in Fig. 17 (b)). Fig. 17 (c) illustrates the detailed multi-AMR navigation system. Due to the limited workspace, the velocity limit in (1) is set to be $u_{\max}=0.2$m/s, where the sensing and collision radii are selected to be $R=1.5$m and $r=1$m by A5. The weights $b,c$ in (19) are set to be $b=10000,c=100$, where the penalty parameter is set to be $\varpi=1000$. The design of the distributed estimators $\widehat{v}_{i}$ for $v_{d}$ in (19) is similar to Section VI-C [41].
@@ VII-B Limitations and Future Work
Despite our effort and exploration for the MPCM navigation in obstacle environments, our work still has some limitations, which require further investigation in future work. Particularly, (i) The proposed CATE algorithm is essentially an MIQP problem, where the computational complexity increases exponentially with the number of robots, making it unsuitable for large-scale scenarios. Future work will consider the fully distributed algorithm for MPCM navigation in obstacle environments to reduce the computational burden. (ii) This paper only considers the simplest smooth circle to cover the irregular obstacles for the usage of CBF constraints, which may sacrifice additional space. Future work will consider more efficient yet non-smooth obstacle boundaries, such as polygons to save space. (iii) This paper only considers the simple integrator dynamics for robots, which is not practical in real applications. Future work will further investigate the complicated underactuated robot dynamics. (iv) This work only considers the scenario of static and moving obstacles at known velocities. Future work will consider the complex MPCM navigation with obstacles of random velocities. (iv) This paper fails to calculate the problem hardness of MPCM navigation due to the integer and real variables. Future work will investigate the problem of hardness with a formal proof.

### v2

@@ V-C Complexity Discussion of Problem 1
According to the property P2 in Definition 2, Problem 1 falls within the category of unlabeled multi-robot motion planning. Based on the summary and comparison of the problem hardness in Table I, it is challenging to determine the hardness of Problem 1. (i) On one hand, if we ignore the robot dynamics, the optimization objective of path-crossing minimization, the well-separation condition of robots, desired points, and obstacles, and only consider static obstacles and 2D environments, there exists a polynomial-time reduction from the special case in [50] to the simplified version of Problem 1, proving that the simplified Problem 1 is also P-SPACE hard. Unfortunately, there is no existing work that proves the theoretical hardness of unlabeled multi-robot motion planning with robot dynamics, although practical solutions have been proposed using efficient sample-based algorithms [57]. If any of the ignored elements are reintroduced, the hardness of Problem 1 may become even harder than the conjectured P-SPACE hardness. (ii) On the other hand, if the radius of the robot is selected to be 4/r, Assumptions A2-A4 in Problem 1 are similar to the well-separation condition in [56], with the slight distinction in the initial relative positions between robots and desired points. This similarity may reduce the hardness of Problem 1 to be solvable in polynomial time. However, as shown in Table I, Problem 1 still involves additional constraints of the robot dynamics, moving obstacles, different optimization objectives, and 3D environments. These factors make it challenging to find a provable and optimal solution that runs in polynomial time. Although we provide an MIQP solution (i.e., the CATE algorithm (19)), this approach cannot be used to determine the hardness of Problem 1 in theory. (iii) Finally, due to the absence of a suitable reduction framework, determining the hardness of Problem 1 under the aforementioned additional constraints is a non-trivial task. This challenge represents a promising direction for future research in its own right.
@@ VI-A Accommodating the CATE to Unicycle Dynamics
TABLE III: Second simulation: Quantitative evaluation metrics of the multi-robot navigation efficiency with 16 different robot-obstacle group settings.
@@
Figure 11: Nine illustrative examples of the “Arrow” formation using the proposed CATE algorithm (19) in the second simulation. In the detailed nine trials, the numbers of robots and obstacles are set to be the same, i.e., $N=11$ and $M=7$, and the size of the obstacle circle is randomly selected in the region of $r\in[1.7,4]$. The positions of the robots and obstacles are randomly selected within the space of $\{x\in[-5,~35],y\in[0,~25]\}$, but do not overlap each other or in the space of $\{x\in[10,~25],y\in[7,~18]\}$. The abbreviations, blue and yellow triangles, red circles, and dashed colored lines have the same meanings as those in Fig. 6.
@@ VI-C Simulations
In the second simulations (see Tables II-III and Figs. 7-12), to better compare the performance with alternative approaches in Section VI-B, we conduct extensive experiments of different robot-obstacle combinations and analyze the aggregated results. Precisely, as shown in Table II, we select 16 groups of different numbers of robots and obstacles (i.e., robots: $N=5,7,9,11$, obstacles $M=4,5,6,7$), and then conduct 10 trials for each group. In each trial, the size of the obstacle circle is randomly selected in the range of $r\in[1.7,4]$. The positions of the robots and obstacles are randomly chosen within the area of $\{x\in[-5,~35],y\in[0,~25]\}$, but do not overlap other robots, obstacles, or in the center space of $\{x\in[10,~25],y\in[7,~18]\}$.
Figure 12: Three failure examples using the previous three different baseline methods in the second simulation. (a) The deadlock failure in the 77-robot-77-obstacle scenario when using the FOTE method. (b) The infeasible failure in the 99-robot-66-obstacle scenario when using the DAPT+ORCA method. (c) The infeasible failure in the $11$-robot-77-obstacle scenario when using the CAPT+ORCA method.
Figure 13: Third simulation: Effectiveness of the CATE algorithm (19) in environments with moving obstacles. (a) T=$1.2$s. (b) T=$3.5$s. (c) T=2020s. The light blue circles represent the initial positions of the obstacles. The abbreviation, blue and yellow triangles, red circles, and dashed colored lines have the same meanings as those in Fig. 6.
@@
As shown on the left side of Table II and Fig. 7, the success rates of the proposed CATE algorithm (19) are $100\%$ all along, which verifies the feasibility advantages of the algorithm in the obstacle environments. The right side of Table II and Fig. 8 illustrates the mean (SD) and boxplots of the convergence time of the four approaches, where the convergence time of the proposed CATE algorithm (19) is the smallest among the ones in three baseline benchmarks. To quantitatively evaluate the multi-robot navigation efficiency, we further employ two metrics for path crossings and trajectory lengths in Table III. It is observed on the left side of Table III that if the number of obstacles and robots increases, the proposed CATE algorithm (19) not only guarantees fewer path crossings than the three baseline methods but also maintains $100\%$ success rates all along. As shown on the right side of Table III, the proposed CATE algorithm (19) only requires the smallest trajectory length than the ones in three baseline methods, which thus indicates the highest efficiency of the CATE algorithm as well. Moreover, Figs. 9 and 10 present the detailed boxplots of the path-crossing and trajectory-length metrics across four approaches, demonstrating that the CATE algorithm (19) achieves the highest performance.
@@
Since the initial positions of robots and obstacles are randomly selected in each trial for each algorithm, one has that these positions are different at each trial. Meanwhile, as shown in Table II, there are considerable differences in the success rates between the CAPT+ORCA and the CATE algorithms (19), whereas we only count the path crossings in Table III when the MPCM navigation is successfully achieved. Therefore, although the CAPT+ORCA algorithm achieves fewer path crossings than the proposed CATE algorithm (19) in Table III, it is not suitable to compare such path crossings directly. Here, we provide a detailed explanation below.
Figure 16: (a) Multi-AMR working space consisting of three different obstacles in the Robotic Research Center (RRC). (b) Two AMRs SCOUTMINI in the experiments and components. (c) 3D point map of the working space in (a) built by a popular LiDAR SLAM framework: Fast-LIO. (d) 2D octomap from the x-y plane projection of the original 3D octomap [63]. (e) Structure of the multi-AMR onboard navigation system, which consists of onboard SLAM, onboard navigation, and onboard execution, respectively.
@@
(i) When the number of robots and obstacles is relatively small (i.e., $N=5,7$ and $M=4,5$), the random selection of their positions results in a high probability that obstacles will not be positioned between the robots and the desired points. Then, such random placement causes the environment to degenerate into an almost obstacle-free one, which is more suitable for the CAPT+ORCA algorithm (this also explains why the success rate of the CAPT+ORCA algorithms is low since it is effective in simplistic environments). (ii) When the numbers of robots and obstacles become relatively large (i.e., $N=9,11$ and $M=6,7$), although the CAPT+ORCA algorithm achieves relatively few path crossings, its success rate is quite low (only 40%-70%, as indicated in Table II). In contrast, the proposed CATE algorithm (19) not only guarantees a 100% success rate (see Table II) but also maintains relatively low path crossings. However, there is a tradeoff between the success rate and the number of path crossings. To achieve a 100% success rate, the CATE algorithm 19 needs to successfully deal with all complicated environments where obstacles effectively obstruct the motion from the initial positions to the desired positions of robots. This results in slightly more path crossings than the CAPT+ORCA algorithm. It’s worth mentioning that the CATE algorithm (19) significantly outperforms the other three algorithms in terms of convergence time and trajectory length in Tables II and III. Therefore, the superiority of the CATE algorithm (19) is verified as well.
@@
Figure 17: First experiment: Comparing platoon formations forming and snapshots using different methods in Robotic Research Center, NTU. (a) FOTE. (b) DAPT + ORCA. (c) CAPT + ORCA. (d) Proposed CATE. The abbreviations, blue, magenta, and yellow triangles, red circles, dashed colored lines have the same meanings as those in Figs. 6 and 15.
@@
In the fifth simulation (see Fig. 15), to show the feasibility of the CATE algorithm (19) for possible deadlocks, we compare our algorithm with the fixed-ordering strategy in a special case of eight robots passing through a narrow gap. In particular, we use two big disk obstacles centered at $\mathbb{x}_{1}^{o}=[6,2]^{\mbox{\tiny\sf T}},\mathbb{x}_{2}^{o}=[6,14]^{\mbox{\tiny\sf T}}$ respectively with the same radius $r_{1}^{o}=r_{2}^{o}=5.5$ to occupy most of the passing space and leave a narrow gap deliberately. The initial positions of robots and desired points are set satisfying A2, A3, A4. As shown in Fig. 15 (a), even though eight robots governed by the CATE algorithm (19) are congestive at the entrance of the narrow gap (magenta vehicles), they can get rid of the congestion and finally, form a platoon (yellow vehicles) by flexibly reallocating their desired points with distinct ordering sequences, which verifies the strong feasibility of the CATE algorithm (19). However, one observes in Fig. 15 (b) that eight robots governed by the fixed-ordering methods fail to pass the gap and are trapped in deadlocks (magenta vehicles). Moreover, Fig. 15 (c) illustrates the evolution of the reallocated ordering sequences $s[i],i\in\mathbb{Z}_{1}^{8},$ for robot $i,i\in\mathcal{V},$ satisfying $s[i]=k,\alpha_{i,k}=1$, which converges to s[1]=4,s[2]=5,s[3]=6,s[4]=2,s[5]=1,s[6]=3,s[7]=8,s[8]=7s[1]=4,s[2]=5,s[3]=6,s[4]=2,s[5]=1,s[6]=3,s[7]=8,s[8]=7 such that $\lim_{t\rightarrow\infty}\mathbb{x}_{i}(t)-\mathbb{x}_{s[i]}^{d}(t)=0$. Note that the final subscripts $s[i],i\in\mathbb{Z}_{1}^{8}$ are consistent with ones in Fig. 15 (a).
@@ VI-D Experiments with Two AMRs
In this subsection, we employ two AMRs (SCOUTMINI [^1] SCOUTMINI: https://global.agilex.ai/products/scout-mini) to validate Theorem 1. As shown in Fig. 16 (a), we establish a multi-AMR workspace, which consists of three different obstacles of a static vehicle, a workbench, and six cones occupying the way. As shown in Fig. 16 (b), each AMR is 0.6150.615m in length and 0.580.58m in width and equipped with an onboard computer: NVIDIA Jetson Xavier NX3222NVIDIA Jetson Xavier NX3: https://www.nvidia.com/en-us/autonomous-machines/embedded-systems/jetson-xavier-nx/ and a Livox Lidar: MID-360333Livox Lidar MID-360: https://www.livoxtech.com/mid-360 or AVIA [^4] Livox Lidar AVIA: https://www.livoxtech.com/avia. However, since the AMR is not equipped with any localization equipment (such as an on-board GPS or off-board motion capture cameras), we utilize the Lidar with the popular FAST-LIO SLAM algorithm [64] to build a 3D point map (see Fig. 16 (c)) and localize itself. Then, we get a 2D map from the projection of the octomap using octotree [63] (i.e., red circles in Fig. 16 (d)). Fig. 16 (e) illustrates the detailed multi-AMR navigation system. Due to the limited workspace, the velocity limit in (1) is set to be $u_{\max}=0.2$m/s, where the sensing and collision radii are selected to be R=1.5R=1.5m and r=1r=1m by A5. The weights $b,c$ in (19) are set to be $b=10000,c=100$, where the penalty parameter is set to be $\varpi=1000$. The design of the distributed estimators $\widehat{v}_{i}$ for $v_{d}$ in (19) is similar to Section VI-C [41].
@@ VII-B Limitations and Future Work
Despite our effort and exploration for the MPCM navigation in obstacle environments, our work still has some limitations, which require further investigation. Particularly, (i) The proposed CATE algorithm is essentially an MIQP solution, where the computational cost increases exponentially with the number of robots, making it unsuitable for large-scale scenarios. Future work will consider the fully distributed algorithm and include a theoretical analysis. (ii) This paper only considers the simplest smooth circle to cover the irregular obstacles for the usage of CBF constraints, which may sacrifice additional space. Future work will consider more efficient yet non-smooth obstacle boundaries, such as polygons. (iii) This paper only considers the simple integrator dynamics for robots, which are not practical in real applications. Future work will further investigate the complicated underactuated robot dynamics. (iv) This work only considers the scenario of static and moving obstacles at known velocities. Future work will consider the complex MPCM navigation with obstacles of random velocities. (iv) The proposed CATE algorithm cannot be utilized to directly determine the hardness of Problem 1 under the additional constraints of robot dynamics, moving obstacles, varying optimization objectives, and 3D environments. Future work will investigate the formal analysis of the problem’s hardness that accounts for these additional constraints.
Acknowledgement
We would like to thank Prof. Kiril Solovey from Israel Institute of Technology for the valuable discussions on the hardness of the problem in unlabeled multi-robot motion planning.